I do not understand this at all. This is not my best subject and I need help please Using the Discriminant Test, what would be your conclusion about the number and nature of the roots of this quadratic equation?
3x2 - 5x – 2 = 0

Answer by lwsshak3(11628)

I do not understand this at all. This is not my best subject and I need help please Using the Discriminant Test, what would be your conclusion about the number and nature of the roots of this quadratic equation?
3x2 - 5x – 2 = 0
**
For given equation:
a=3, b=-5, c=-2
discriminant=b^2-4ac=25-4*3*-2=25+24=49
In this case, the discriminant is a non=zero positive perfect square, therefore, the solution are two unequal real numbers, rational or integers. For given problem, x=-b±√49=5±7=12 or -2
..
If the discriminant is a positive nonperfect square, the solution would be two unequal irrational numbers.
..
If the discriminant is negative, the solution would be nonreal numbers, called imaginary numbers.